OpenStudy (anonymous):

fan/medal A father in his will left all his money to his children in the following manner: $1000 to the first born and 1/10 of what then remains, then $2000 to the second born and 1/10 of what then remains, then $3000 to the third born and 1/10 of what then remains, and so on. When this was done each child had the same amount. How many children were there?

OpenStudy (anonymous):

It asks how many children are there 81000%2F9000 = 9 children each getting 9000

OpenStudy (mitu12):

n = n - 1 + (2 n - 1 n + 1) / 10 (dividing by 1000) 0 = -1 + (n+1) / 10 n+1 = 10 n = 9 there are 9 children